What is the definition for the "root of the equation" when doing algebra?
Thanks
THE ZERO OF THE QUADRATIC FUNCTION, THAT IS, VALUES OF X WHICH F(X)=0, ARE THE ROOTS OF THE EQUATION.

Found 2 solutions by AnlytcPhil, 55^10:
Answer by AnlytcPhil(1810)

What is the definition for the "root of the equation" when doing algebra?
Thanks
"Root" is a more formal way to say "answer" or "solution". A number is
called a root of an equation if when the number is substituted into the
equation and both sides simplified, the result is an identity, such as
2=2 or 8=8, etc.
Example:
There are two roots to this equation
x²+6=5x
They are 3 and 2. 3 is one root because when you substitute 3 for x
3²+6=5(3)
9+6=15
15=15
Also 2 is a root, because when you substitute 2 for x
2²+6=5(2)
4+6=10
10=10
Edwin

Answer by 55^10(2)